Adult tickets for the game cost $4.00 each and student tickets cost $2.00 each. A total of 920 tickets worth $2,446.00 were sold. How many student tickets were sold?
Answer by mananth(16946)

Adult ticket x Numbers
Child ticket ------------ y Numbers
x+y=920 .............1
4x+2y=2446 .............2

multiply (1)by -2
Multiply (2) by 1
-2x-2y= -1840
4x+2y=2446
Add the two equations
2x=606
/2
x=303
plug value of x in (1)
x+y=920
303+y=920
y=920-303
y=617
y=617